OpenStudy (anonymous):

1)Fred needs 55 ml of 23% sucrose solution to do his science experiment. He needs to mix a 20% sucrose solution with some 30% sucrose solution to get the 23% solution he wants. How much of each should Fred use? (Include units with your numerical answers.) Fred should use________of 20% sucrose solution and ________of 30% sucrose solution. 2) A candy store owner has chocolate candies worth $2.00 per pound and sour candies worth $0.80 per pound. How much of each kind of candies should she combine to get 70 lb of mixed candies worth $1.40 per pound?

OpenStudy (phi):

amount of sucrose is volume * % the final solution will have 55 ml * 0.23 if you start with X ml of the weak solution, it will have X * 0.2 of sucrose if you have Y ml of the strong solution, it will have Y * 0.3 of sucrose you want X+Y= 55 ml (so that you end up with 55 ml of the final solution) you want the total amount of sucrose to be equal: 0.2x + 0.3y = 0.23*55 you have two equations, and two unknowns. Can you solve them ?